OpenStudy (anonymous):

number to complete the square for 3x^2+x-4

OpenStudy (tkhunny):

\(3x^{2} + x - 4\) \(3x^{2} + x) - 4\) \(3(x^{2} + (1/3)x + ______) - 4 - 3(________)\) (1/3)/2 = 1/6 (1/6)^2 = 1/36 \(3(x^{2} + (1/3)x + 1/36) - 4 - 3(1/36)\)